# ReZero: Boosting MCTS-based Algorithms by Just-in-Time and Speedy Reanalyze

### v1

@@ 1 Introduction
Within this field, Monte Carlo Tree Search (Świechowski et al., 2023) (MCTS) has been proven to be a efficient method for utilizing models for planning. It incorporates the UCB1 algorithm (Auer et al., 2002) into the tree search process and has achieved promising results in in a wide range of application scenarios. Specifically, AlphaZero (Silver et al., 2017) plays a big role in combining deep RL with MCTS, achieving notable accomplishments that can beat top-level human players. While AlphaZero can only be applied to environments with perfect simulators, MuZero (Schrittwieser et al., 2019) extended the algorithm to cases without known environment models, resulting in good performances in a wider range of tasks. Following MuZero, many successor algorithms have emerged, enabling MuZero to be applied in continuous action spaces (Hubert et al., 2021), environments with stochastic dynamics (Antonoglou et al., 2021), offline RL training scenarios (Schrittwieser et al., 2021), and etc. All these MCTS-based algorithms (Silver et al., 2016, 2017; Schrittwieser et al., 2019; Ye et al., 2021) made valuable contributions to the universal applicability of the MCTS+RL framework.
@@ 2.1 MCTS-based Algorithms
AlphaGo (Silver et al., 2016) and AlphaZero (Silver et al., 2017) combined MCTS with deep reinforcement learning, achieving significant results in board games, repeatedly defeating world-champion players, and revealing their remarkable capabilities. Following these achievements, MuZero (Schrittwieser et al., 2019) algorithm combines tree search with a learned value equivalent model (Grimm et al., 2020b), successfully promoting the algorithm’s application to scenarios without known models, such as Atari. Subsequent to MuZero’s development, several new works based on the MuZero framework emerged. Ye et al. (2021) further improved MuZero’s sample efficiency, using very little data for training and still achieving outstanding performance. Hubert et al. (2021) and Antonoglou et al. (2021) extended the MuZero agent to environments with large action spaces and stochastic environments, further broadening the algorithm’s usage scenario. Schrittwieser et al. (2021) applied MuZero’s reanalyze operation as a policy enhancement operator and extended it to offline settings. Danihelka et al. (2022) changed the node selection method during tree search, so the policy obtained from the search shows a monotonic improvement, ensuring algorithm performance even with low simulation numbers. Despite the aforementioned excellent improvement techniques, the current time overhead of MCTS-based RL remains significant, which is the main problem this paper aims to address.
@@ 4.2.2 Theoretical Analysis
The upper bound given by Theorem 4.3 is $O(\sqrt{n})$, which, while not reaching the level of $O(\log{n})$, is still sub-linear. Inequality (16) also indicates that a better policy which has higher probability to choose the best arm can lower the upper bound. Combining 4.2 and 4.3, we can conclude that with the increase in the number of simulations, the visit distribution will increasingly concentrate on the optimal arm, and the value corresponding to the policy will also converge to the value of the optimal arm. If the assumption that $P_{1}>\mu_{1}$ doesn’t hold, an upper bound with $O(\sqrt{n})$ rate can still be obtained by similar proof as long as there exists a $t$ that $P_{1}\sqrt{t-1}>\mu_{1}$, which indicates unlike some algorithms that are sensitive to prior policies, MCTS has strong robustness to prior policies. As long as the number of simulations is sufficient, it can still achieve a good target policy.
@@
Table 1: Average wall-time of ReZero-M vs. MuZero on various tasks. (left) Six Atari games, (right) two board games. The time represents the average total wall-time to 100k environment steps for each algorithm. Mean and standard deviation over 5 runs. ReZero used two to four times less wall-time per 100k steps than the baseline method.

### v2

@@ 1 Introduction
Within this field, Monte Carlo Tree Search (Świechowski et al., 2023) (MCTS) has been proven to be a efficient method for utilizing models for planning. It incorporates the UCB1 algorithm (Auer et al., 2002) into the tree search process and has achieved promising results in in a wide range of application scenarios. Specifically, AlphaZero (Silver et al., 2017) plays a big role in combining deep RL with MCTS, achieving notable accomplishments that can beat top-level human players. While AlphaZero can only be applied to environments with perfect simulators, MuZero (Schrittwieser et al., 2019) extended the algorithm to cases without known environment models, resulting in good performances in a wider range of tasks. Following MuZero, many successor algorithms have emerged, enabling MuZero to be applied in continuous action spaces (Hubert et al., 2021), environments with stochastic dynamics (Antonoglou et al., 2021), offline RL training scenarios (Schrittwieser et al., 2021), and etc. All these MCTS-based algorithms (Silver et al., 2016, 2017; Schrittwieser et al., 2019) made valuable contributions to the universal applicability of the MCTS+RL framework.
@@ 2.1 MCTS-based Algorithms
AlphaGo (Silver et al., 2016) and AlphaZero (Silver et al., 2017) combined MCTS with deep RL, achieving significant results in board games, defeating world-champion players, and revealing their remarkable capabilities. Following these achievements, MuZero (Schrittwieser et al., 2019) combines tree search with a learned value equivalent model (Grimm et al., 2020b), successfully promoting the algorithm’s application to scenarios without known models, such as Atari. Subsequent to MuZero’s development, several new works based on the MuZero framework emerged. Ye et al. (2021) further improved MuZero’s sample efficiency, using very little data for training and still achieving outstanding performance. Hubert et al. (2021) and Antonoglou et al. (2021) extended the MuZero agent to environments with large action spaces and stochastic environments, further broadening the algorithm’s usage scenario. Schrittwieser et al. (2021) applied MuZero’s reanalyze operation as a policy enhancement operator and extended it to offline settings. Danihelka et al. (2022) changed the node selection method during tree search, so the policy obtained from the search shows a monotonic improvement, ensuring algorithm performance even with low simulation numbers. Despite the aforementioned excellent improvement techniques, the current time overhead of MCTS-based RL remains significant, which is the main problem this paper aims to address.
@@ 4.2.2 Theoretical Analysis
The upper bound given by Theorem 4.3 is $O(\sqrt{n})$, which, while not reaching the level of $O(\log{n})$, is still sub-linear. Inequality (16) also indicates that a better policy which has higher probability to choose the best arm can lower the upper bound. Combining 4.2 and 4.3, we can conclude that with the increase in the number of simulations, the visit distribution will increasingly concentrate on the optimal arm, and the value corresponding to the policy will also converge to the value of the optimal arm. If the assumption that $P_{1}>\mu_{1}$ doesn’t hold, an upper bound with $O(\sqrt{n})$ rate can still be obtained by similar proof as long as there exists a $t$ that $P_{1}\sqrt{t-1}>\mu_{1}$, which indicates unlike some algorithms that are sensitive to prior policies, MCTS has strong robustness to priors. As long as the number of simulations is sufficient, it can still achieve a good target policy.
@@
Table 1: Average wall-time of ReZero-M vs. MuZero on various tasks. (left) Six Atari games, (right) two board games. The time represents the average total wall-time to 100k environment steps for each algorithm. Mean and standard deviation over 5 runs.
@@ 7 Impact Statements
8 Acknowledgements
Thanks to each of the authors for their efforts in the work of ReZero. We spent several memorable mid-nights and early mornings together in the lab. Thanks to everyone at OpenDILab and their pioneering work, especially the LightZero benchmark (https://github.com/opendilab/LightZero) for providing the base code for ReZero. Finally, we would like to thank Shanghai AI Laboratory and professor Jing Yang of Xi’an Jiaotong University for providing a platform for this work.